how to calculate enterprise value if we have free cash flow and will remain constant forever and WACC
Answer #1

When free cash flow will remain constant forever then it is in the form of a perpetuity. The value of a perpetuity = constant annual amount/discount rate. Here the free cash flow amount is the constant annual amount and WACC is the discount rate.

Supposing that the company has no debt then enterprise value = Amount of constant free cash flow/WACC. For instance if the amount of free cash flow per year is $10,000 and WACC is 8% then enterprise value = $10,000/8% = $125,000

If the company has debt then use the free cash flow figure to calculate the unlevered free cash flow. It is also known as the free cash flow to the firm. Suppose that a company has an amount of unlevered free cash flow of $15,000 and this amount will remain constant in future. Its WACC is 9%. So its enterprise value = 15000/9% = $166,666.67